Circular No. 5944

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1994 R. M. Wagner, Lowell Observatory, reports that a spectrum obtained

on Mar. 4 by R. Bertram shows an “un-novalike” spectrum with a very red continuum and what may be narrow Hβ emission. There are also many narrow absorption lines and a broader absorption at ∼ 500 nm.

P. M. Kilmartin gave the 2000.0 version of the position on IAUC 5943 as α = 18^{h}50^{m}36^{s}.65, δ = −21^{o}23′28′′.5. Her FK_{5}/J2000.0 re-reduction using the PPM gives end figures 36^{s}.73, 29′′.1 and 36^{s}.74, 28′′.4 for the separate exposures. B. A. Skiff, Lowell Observatory, provides the following end figures for a PPM reduction of his measurements of a star on the Palomar Sky Survey prints close to this position: 36^{s}.83, 29′′.5 (R = 16), 36^{s}.77, 29′′.1 (B = 18). There is a fainter star (R = 19) at end figures 36^{s}.49, 27′′.8.

Further photometry by A. C. Gilmore with the 0.6-m f/16 Cassegrain telescope at Mt. John University Observatory: Mar. 3.67 UT, V = 8.67, U − B = +1.22, B − V = +1.65, V − R = +0.93, V − I = +1.82 (marginal conditions); 4.67, 8.76, +1.4, +1.73, +0.96, +1.91 (good conditions). The general uncertainty of ±0.05 mag (0.1 mag for U − B on Mar. 4) appears also to be valid for the Feb. 27.68 observation (IAUC 5943).

Visual magnitude estimates by F. Farrell, Hackham, South Australia: Feb. 28.77 UT, 7.9; Mar. 3.79, 8.8.

GRB 940301 K. Hurley, University of California; G. Fishman, C. Kouveliotou and C.

Meegan, NASA/Marshall Space Flight Center; T. Cline, NASA/Goddard Space Flight Center; and M. Boer and M. Niel, Centre d’Etude Spatiale des Rayonnements, Toulouse, report that the annulus of arrival of the γ-ray burst detected by Ulysses, BATSE, and COMPTEL (IAUC 5943) on Mar 1.84071 UT is centered at α = 21^{h}27^{m}.65, δ = +52^{o}30′.46 (equinox 2000.0) with radius 59^{o}.26 and total width 0^{o}.30. Further analysis will reduce this width, but the present annulus is fully compatible with the COMPTEL and BATSE error regions, as well as with the EGRET error region.

OJ 287 Visual magnitude estimates (viz. the prediction on IAUC 5909): Mar.

2.85 UT, [14.9 (T. Vanmunster, Landen, Belgium); 3.85, 14.5 (Vanmunster); 3.87, 14.6 (P. Van Cauteren, Aartselaar, Belgium); 3.99, 14.6 (Vanmunster); 4.89, 14.7 (Vanmunster).

1994 March 5 Brian G. Marsden